OpenStudy (anonymous):

Use the binomial coefficient formula to calculate the binomial coefficient for the 7th term in a polynomial expansion when the power of the binomial is 8.

OpenStudy (anonymous):

\[\left(\begin{matrix}8 \\ 7\end{matrix}\right)x^1y^7= 8x^1y^7\]